Grade 11Physical Chemistry

You are given one litre of 0.15M HCl and one litre of 0.40 M HCl. What is the maximum volume of 0.25 HCl which you can make from these solutions without adding any water?

Abhishek Kumar
8 Years agoGrade 11

2 Answers

Om
8 Years ago
Given 
one litre of 0.15M HCl and one litre of 0.40 M HCl 
so there would be a total of 0.55 moles and 2L volume with molarity 0.275
if you only want molarity as 0.25 then maximum volume would be 1.81L
 
 
NIRMALYA CHAKRABARTY
5 Years ago
Let x lit of 0.40M HCl added to 1 lit of 0.15M HCl. 
Total Vol. = (1+x)
M1V1 + M2V2 = M3V3 
0.4x + 0.15*1 = 0.25(1+x)
x = 0.67 L
So total Vol. Of HCl  solution is (1+0.67)= 1.67 L